OpenStudy (anonymous):

1. find the limit x approaches 0, f(x)=x/|x| 2. find the limit x approaches 2, f(x)=|x-2|/x^2+x-6

OpenStudy (anonymous):

\[|x|=\begin{cases} x&\text{for }x\ge0\\-x&\text{for }x<0 \end{cases}\] which means \[|x-2|=\begin{cases} x-2&\text{for }x\ge0\\-(x-2)&\text{for }x<0 \end{cases}\] So for the first limit, when \(x\to0\), consider what happens when \(x\to0\) from either side. That is, consider the one-sided limits \[\lim_{x\to0^-}f(x)~~~~\text{and}~~~~\lim_{x\to0^+}f(x)\] If \(x\to0^-\) (approaching from the left), then that means any value of \(x\) would be negative, i.e. \(x<0\). This would mean \(|x|=-x\), so \(f(x)=\dfrac{|x|}{x}=\dfrac{-x}{x}=-1\). The limit would then be -1. From the other side, if \(x\to0^+\), then \(x>0\), so \(|x|=x\) and \(f(x)=\dfrac{|x|}{x}=\dfrac{x}{x}=1\), and so the limit is 1. The one-sided limits are not equivalent, so \(\displaystyle\lim_{x\to0}\frac{|x|}{x}\) does not exist.

OpenStudy (anonymous):

Use this as an example to reason through the second question.

OpenStudy (anonymous):

in the second question , the absolute value is the numerator. can i still use the same approach?

OpenStudy (anonymous):

You can. The denominator can be factored.

OpenStudy (anonymous):

\[\frac{ \left| x -2 \right| }{ \left( x +3 \right)\left( x -2 \right) }\] \[=\frac{ 1 }{ x +3 }\] from the left,\[\frac{ 1 }{ -\left( 2+3 \right) }=\frac {- 1}{5}\] and the right will be \[\frac{ 1 }{ 5 }\] is that it?